Circular No. 5996

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

MARKARIAN 421 A. D. Kerrick reports for the Whipple Observatory Gamma Ray Col-

laboration (Iowa State University, University of Michigan, Purdue University, Smithsonian Astrophysical Observatory, University of Leeds, and the National University of Ireland): “Observations of Mkn 421 with the Whipple Observatory 10-m γ-ray reflector on May 15 showed a significant increase in the TeV γ-ray emission. Over the period 1994 Jan.–Apr., the TeV γ-ray flux was detected at a level that was a factor of ∼ 2 below that previously observed (Punch et al. 1992, Nature 358, 477). As late as May 10, the flux was consistent with this low level; on May 14, it was observed to increase, reaching a level on May 15 of ten times that of the Jan.–Apr. flux. The observed detection rate was 3.9 counts/minute. At the energy threshold of 250 GeV, this corresponds to a flux of 1.87 ± 0.34 × 10^{−10} photon cm^{−2} s^{−1}, which is comparable to the flux from the Crab Nebula at this energy. No observations were possible after May 15 because of the moon. The x-ray flux was reported at a high level on May 16 (IAUC 5993).”

COMET TAKAMIZAWA (1994i) Improved parabolic orbital elements from MPEC 1994-K04:

T = 1994 June 29.5773 TT  ω = 192.^{o}1029
Ω = 51.1232 }2000.0
q = 1.948146 AU  i = 135.9779
1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 23 15^{h}21.^{m}29 − 8^{o}41.^{′}6 1.007 2.002 165.^{o}2 7.^{o}4 10.0
28 14 45.86 − 9 50.1 1.024 1.989 154.9 12.5 10.0
June 2 14 12.82 −10 46.3 1.070 1.978 143.2 17.9 10.1
7 13 43.64 −11 30.2 1.138 1.968 132.1 22.5 10.2
12 13 18.89 −12 04.4 1.225 1.960 121.8 26.1 10.4
17 12 58.45 −12 32.0 1.325 1.954 112.5 28.7 10.5
22 12 41.83 −12 56.0 1.435 1.950 104.1 30.4 10.7
27 12 28.43 −13 18.3 1.551 1.948  96.5 31.2 10.8
July 2 12 17.68 −13 40.3 1.670 1.948  89.6 31.5 11.0
7 12 09.07 −14 02.8 1.790 1.950  83.1 31.2 11.2
12 12 02.20 −14 26.3 1.910 1.954  77.2 30.5 11.3
17 11 56.75 −14 51.2 2.027 1.960  71.6 29.5 11.5

1994 May 25 Daniel W. E. Green